OpenStudy (anonymous):

is my answer right? my answer is 19.2 i just need to get this question right and ill graduate on time plzz help.

OpenStudy (anonymous):

There no question

OpenStudy (anonymous):

Yeah....

OpenStudy (anonymous):

what is the question?

OpenStudy (anonymous):

we cant know if you're right if there is no question

OpenStudy (anonymous):

Lol Im lost haha

OpenStudy (anonymous):

im sorry i took for ever my computer is lagging lol

OpenStudy (anonymous):

Its almost identical to the last question we did

OpenStudy (anonymous):

19.2 is the area of 2 triangle only

OpenStudy (anonymous):

dont forget the square

OpenStudy (whpalmer4):

I can't see the whole thing, but aren't you supposed to find the area of only triangle AEF?

OpenStudy (whpalmer4):

height is 6, base is 3.2, area of a triangle is \[A = \frac{1}{2}b*h\]

OpenStudy (anonymous):

is it 8.6?

OpenStudy (whpalmer4):

If \[\frac{1}{2}*6*3.2=8.6\]then yes. If not, then no.

OpenStudy (anonymous):

isnt that just area of triangle?

OpenStudy (whpalmer4):

as I said, I can't see the entire question, but the part I can see appears to be asking for the area of triangle AEF.

OpenStudy (anonymous):

oo ok i thought they wanted the paralellogram

OpenStudy (whpalmer4):

the original poster will have to read it and decide what it is asking

OpenStudy (anonymous):

i just need to find the area of AEF

OpenStudy (anonymous):

Area of AEF = 19.2, yes. You are correct

OpenStudy (anonymous):

You multiply Base * height

OpenStudy (anonymous):

but do i have to take have of 19.2?

OpenStudy (anonymous):

Triangle AEF just use \[\frac{ 1 }{ 2 }\times base \times height\]

OpenStudy (anonymous):

should be just 9.6

OpenStudy (whpalmer4):

At least someone here knows how to find the area :-)

OpenStudy (anonymous):

thank you so much kingdomdream, i need to study more but i pass my test ;)

OpenStudy (anonymous):

no problem =)